person A takes 3 hours to work while person B takes 4 hours. how long does it take them when working together?
Found 2 solutions by fcabanski, mathstutor494:
Answer by fcabanski(1391)

Problems of this type come down to finding each person's work rate, adding the rates, then determining how long to complete the entire job.


A completes the whole job in 3 hours, and so completes 1/3 of the job per hour.
B completes the whole job in 4 hours, and so completes 1/4 of the job per hour.
Working together they complete 1/3 + 1/4 = 7/12 of the job per hour.


1 (the whole job) = rate * time. The rate is 7/12 job per hour, so:
1=7/12 + t


Divide both sides by 7/12:
1*12/7 = t (in hours)
12/7 = t (in hours) = approximately 1.7 hours = 1 hour 42 minutes.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com


Answer by mathstutor494(120)

Person A takes 3 hours to work, so he completes 1/3 work in 1 hr.
Similarly,Person B takes 4 hours to work, so he completes 1/4 work in 1 hr.

Working togther they complete 1/3+ 1/4 = 7/12 in 1 hr.
Hence total time taken to complete work = 12/7 hrs (Unitory Method)